# An equivalence in random matrix and tensor models via a dually weighted intermediate field representation

### v1

@@ Discussion.
The subsection 4.3 is the core of the Section 4: the complex matrix model (4.1) and the self-adjoint matrix model (4.7) with the logarithmic intermediate potential (4.14) are not merely identified as the same power series—they are exactly the same theory once expressed in the variables. The field $B$ plays the role of an intermediate Fourier variable that enforces $A=M^{\dagger}M$, while its potential $Y[P,Q]$ encodes the PP– and QQ–dependence arising from integrating out $M$. Thus, the two models are dual descriptions of the same underlying matrix integral, and any expectation value depending on $M^{\dagger}M$ may be computed in whichever formulation is more convenient.
@@ 4.4 Examples
$V[K,L](M^{\dagger}M)=-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes 2}-iL\otimes(KM^{\dagger}M)\right)\;.$ (4.28)
Its partition function reads
$\mathcal{Z}_{\rm CM}[V[K,L]](P,Q)=\int_{\mathbb{C}^{N\times N}}{\cal D}M^{\dagger}{\cal D}M\;e^{-N{\rm Tr}(P^{-1}M^{\dagger}Q^{-1}M)-{\rm Tr}\ln({\mathbb{1}}^{\otimes 2}-iL\otimes(KM^{\dagger}M))}\;.$ (4.29)
As in Section 4.1, contractions of the first index of the matrix $M$ generate color–1 faces weighted by ${\rm Tr}(Q^{n})$. Given the form of the dually weighted potential $V[K,L]$, each vertex carries a weight ${\rm Tr}(L^{n})$, just as in the self-adjoint one-matrix model. However, in this case, the face structure is enriched by the presence of two independent families of cycles built into the complex matrix model. This is because contractions of the second index of the matrix $M$ generate color–2 faces weighted by ${\rm Tr}((PK)^{n})$. This is illustrated in Figure 5. In summary, for each $n$, one obtains
@@
$\mathcal{Z}_{\rm HM}[V[K,L],Y[P,Q]]=\int dAdB\;e^{-iN{\rm Tr}(AB)-{\rm Tr}\ln({\mathbb{1}}^{\otimes 2}-iL\otimes(KA))-{\rm Tr}\ln({\mathbb{1}}^{\otimes 2}-iQ\otimes(PB))}\;.$ (4.30)
@@
  • AA-vertices of degree $n$ with weight ${\rm Tr}(L^{n})$.
  • $B$-vertices of degree $n$ with weight ${\rm Tr}(Q^{n})$.
  • Faces of length $n$ with weight ${\rm Tr}((PK)^{n})$.
This is a bipartite-graph version of dually weighted matrix models [Kazakov_1996, das_1990], and according to Theorem 4.3, we have
$\frac{\mathcal{Z}_{\rm CM}[V[K,L]](P,Q)}{\mathcal{Z}_{\rm CM}[0](P,Q)}=\frac{\mathcal{Z}_{\rm HM}[V[K,L],Y[P,Q]]}{\mathcal{Z}_{\rm HM}[0,Y[P,Q]]}\;.$ (4.31)
Figure 5: An example of equivalence between a graph in the complex and the self-adjoint matrix models. Here, the colors red, green, blue and purple represents factors of $P$, $K$, $Q$ and $L$, respectively.
@@
An interesting case arises when $Q=C_{2}$ (See (3.6)). We recall that the matrix $C_{2}$ is only well defined in the large $N$ limit [Benedetti:2008hc]. Therefore, consider the large-$N$ partition function
$\mathcal{Z}_{\rm HM}[V,Y[P,C_{2}]]=\int dAdB\;e^{-iN{\rm Tr}(AB)+\frac{1}{2}N{\rm Tr}((PB)^{2})+V(A)}\;.$ (4.32)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,C_{2}]]=N^{-n}\sum_{\gamma\in S_{n}}{\rm Tr}_{[\gamma\sigma]}(P){\rm Tr}_{[\gamma]}(C_{2})\;.$ (4.34)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,C_{2}]]=0\qquad\text{if $n$ is odd}\;,$ (4.35)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y[P,C_{2}]]=N^{-m}\sum_{\gamma\in[2^{m}]}{\rm Tr}_{[\gamma\sigma]}(P)\qquad\text{if $n=2m$ is even}\;.$ (4.36)
@@
for $n=2m$ even and $1\leq n\leq N$. This correspondence is illustrated in Figure 6 and the explicit example in Figure 7.
Figure 6: An example of equivalence between a graph in the complex model with $Q=C_{2}$ with $C$ theory ${\cal Z}_{\rm CM}[V](P,C_{2})$ (4.37) and the Hermitian matrix theory $\mathcal{Z}_{\rm HM}[V,Y[P,C_{2}]]$ (4.33).
Figure 7: An example of equivalence between a graph in the complex matrix theory ${\cal Z}_{\rm CM}[V](P,C_{2})$ (4.37) and the Hermitian matrix theory $\mathcal{Z}_{\rm HM}[V,Y[P,C_{2}]]$ (4.33). The graph on the right is obtained by gluing (or contracting) the red and green faces of the graph on the left.
@@
where ${\rm Cat}(m)$ are the Catalan numbers. If $V(z)=-N\frac{g^{2}}{2}z^{2}$, then
$\mathcal{Z}_{\rm HM}[V,Y[{\mathds{1}},C_{2}]]=\int dA\;e^{-N{\rm Tr}\left(\frac{1}{2}A^{2}-\frac{g^{2}}{2}A^{2}\right)}\;,$ (4.40)
showing that the model (3.10) becomes equivalent to the Gaussian model above. Appendix E explores the case when $Q=C_{2}$ further via explicit graph resummation.
$\mathbf{Q=C_{k}}$: a $k$-uniform rigidity and a partition-matroid structure.
@@
$\big\langle{\rm Tr}_{[\sigma]}(M^{\dagger}M)\big\rangle[0](P,C_{k})=\big\langle{\rm Tr}_{[\sigma]}(A)\big\rangle[0,Y[P,C_{k}]]=N^{-d}\sum_{\gamma\in[k^{d}]}{\rm Tr}_{[\gamma\sigma]}(P).$ (4.44)
@@ 4.5 Miscellaneous results
Let $A\in\mathrm{Herm}(N)$ be a random matrix distributed according to ${\cal Z}_{\mathrm{HM}}[0,Y[P,Q]]$. Then, for any representation $r$ of ${\rm GL}(N)$,
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=n!N^{-n}\frac{1}{\chi^{r}(e)}\chi_{r}(P)\chi_{r}(Q)\;,$ (4.46)
@@
$\chi_{r}(A)=\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi^{r}(\sigma){\rm Tr}_{[\sigma]}(A)\;.$ (4.47)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=\frac{1}{n!}N^{-n}\sum_{\begin{subarray}{c}\sigma,\gamma\in{\rm S}_{n}\end{subarray}}\chi^{r}(\sigma){\rm Tr}_{[\gamma\sigma]}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (4.48)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=\frac{1}{n!}N^{-n}\sum_{\begin{subarray}{c}\sigma,\gamma\in{\rm S}_{n}\end{subarray}}\chi^{r}(\sigma)\sum_{s}\chi^{s}(\gamma\sigma)\chi_{s}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (4.49)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=\frac{1}{n!}N^{-n}\sum_{\begin{subarray}{c}\sigma,\gamma\in{\rm S}_{n}\end{subarray}}\chi^{r}(\sigma)\sum_{s}\frac{\chi^{s}(\gamma)\chi^{s}(\sigma)}{\chi^{s}(e)}\chi_{s}(P){\rm Tr}_{[\gamma]}(Q)\;,$ (4.50)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=N^{-n}\sum_{\begin{subarray}{c}\gamma\in{\rm S}_{n}\end{subarray}}\sum_{s}\frac{\chi^{s}(\gamma)\delta_{rs}}{\chi^{s}(e)}\chi_{s}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (4.51)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=N^{-n}\sum_{\begin{subarray}{c}\gamma\in{\rm S}_{n}\end{subarray}}\frac{\chi^{r}(\gamma)}{\chi^{r}(e)}\chi_{r}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (4.52)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=n!N^{-n}\frac{1}{\chi^{r}(e)}\chi_{r}(P)\chi_{r}(Q)\;.$ (4.53)
This expression is valid for any $N>0$ and $n\geq 0$. If $n\leq N$, we can use $\chi^{r}(e)=N^{-n}n!\chi_{r}(C_{1})$. For the following corollary, we recall the matrix $C_{1}$ given in (3.6), the partition function ${\cal Z}_{\mathrm{HM}}$ given in (4.7), and the expectation value $\langle\cdot\rangle$ given in (4.8).
@@
Let $A\in\mathrm{Herm}(N)$ be a random matrix distributed according to ${\cal Z}_{\mathrm{HM}}[0,Y[P,Q]]$. Then, for any representation $r$ of ${\rm GL}(N)$ with size $n\leq N$,
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=\frac{\chi_{r}(P)\chi_{r}(Q)}{\chi_{r}(C_{1})}\;.$ (4.54)
Since the character $\chi^{r}(e)$ does not depend on $N$, one has that, at fixed $n$,
$\chi^{r}(e)=n!\lim_{N\rightarrow\infty}N^{-n}\chi_{r}(C_{1})\;.$ (4.55)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y[P,Q]]=\frac{\chi_{r}(P)\chi_{r}(Q)}{\tilde{\chi}_{r}(C_{1})}\;.$ (4.57)
@@
$\left\langle f(A)\right\rangle[0,Y[P,Q]]=\sum_{n=0}^{\infty}\sum_{r\vdash n}f^{r}\frac{\chi_{r}(P)\chi_{r}(Q)}{\tilde{\chi}_{r}(C_{1})}\;.$ (4.59)
@@
$\chi_{r}(M^{\dagger}M)=\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi^{r}(\sigma){\rm Tr}_{[\sigma]}(M^{\dagger}M)\;,$ (4.60)
@@ 5.1 The complex tensor model
If $\varphi$ is a random tensor distributed according to the partition function ${\cal Z}_{\rm CT}[W](R)$ (5.1), then, for any $\bm{\sigma}\in{\rm S}_{n}^{D}$, the $\bm{\sigma}$}–bubble/trace invariant satisfies
@@
$\langle\varphi\varphi^{\dagger}\rangle[0](R)=\bigotimes_{c=1}^{D}R_{c}\;.$ (5.6)
Furthermore, the expectation value (5.4) factorizes accordingly, as stated in the follwoing Corollary.
@@
This factorized expression has a transparent geometric meaning in the stranded Feynman-graph representation of tensor models: it decomposes into $D$ independent families of faces, one for each tensor index. This is because when the covariance factorizes as $R=\bigotimes_{c=1}^{D}R_{c}$, the tensor propagator splits into $D$ independent strands, where the $c$-th strand is proportional to $R_{c}$. This implies that the Wick contractions separate by color: the same permutation $\mu\in S_{n}$ governs the pairings of all tensor slots, but in each color $c$, it combines with the bubble permutation $\sigma_{c}$ to form $\mu\sigma_{c}$. The cycles of $\mu\sigma_{c}$ are precisely the faces traced by the $c$-th index of the tensor, and a cycle of length $l$ corresponds to a color-$c$ face of length $l$ in the stranded graph. Since a color-$c$ strand has a propagator $R_{c}$, a color-$c$ face of length $l$ has a weight ${\rm Tr}(R_{c}^{l})$. This is the tensor analogue of the dually weighted structure of the complex matrix model presented in Section 4.1, now extended to $D$ colors.
@@
In particular, the scalar-$U(1)$ and partial-$U(N)$ cases will be relevant for the equivalence theorems we prove, and we elaborate on these below. Notice that scalar-$U(1)$ $\subset$ partial-$U(N)$ $\subset$ full-$U(N)^{D}$, and that scalar-$U(1)$ $\subset$ repeated-$U(N)$ $\subset$ full-$U(N)^{D}$, but there is no inclusion between partial-$U(N)$ and repeated-$U(N)$.
@@
The difference between a general $W(\varphi,\varphi^{\dagger})$ and a potential of the form $V(\varphi\varphi^{\dagger})$ is illustrated in Figure 8: in the former, $\varphi$ and $\varphi^{\dagger}$ may appear in arbitrary orderings or in different amounts, while in the latter they are naturally paired into copies of $\varphi\varphi^{\dagger}$.
Figure 8: (a) A potential $W$ without scalar-$U(1)$ symmetry. (b) A potential $V$ with scalar-$U(1)$ symmetry but without repeated-$U(N)$ symmetry. (c) A potential $V$ with repeated-$U(N)$ symmetry but without partial-$U(N)$ symmetry. (d) A potential $V$ with partial-$U(N)$ symmetry but without full-$U(N)^{D}$ symmetry. (e) A potential $V$ with full-$U(N)^{D}$ symmetry. Illustration of several polynomial interactions with different types of symmetries.

### v2

@@ Discussion.
The subsection 5.3 is the core of the Section 5: the complex matrix model (5.1) and the self-adjoint matrix model (5.7) with the logarithmic intermediate potential (5.14) are equivalent random geometry theories. The field $B$ plays the role of an intermediate Fourier variable that enforces $A=M^{\dagger}M$, while its potential $Y_{\ln}[P,Q]$ encodes the PP– and QQ–dependence arising from integrating out $M$. Thus, the two models are dual descriptions of the same underlying matrix integral, and any expectation value depending on $M^{\dagger}M$ may be computed in whichever formulation.
@@ 5.4 Examples
$V_{\ln}[K,L](M^{\dagger}M)=-{\rm Tr}\ln\left({\mathbb{1}}^{\otimes 2}-iL\otimes(KM^{\dagger}M)\right)\;$ (5.29)
for constant $N\times N$ matrices $K$ and $L$. Its partition function reads
$\mathcal{Z}_{\rm CM}[V_{\ln}[K,L]](P,Q)=\int_{\mathbb{C}^{N\times N}}{\cal D}M^{\dagger}{\cal D}M\;e^{-N{\rm Tr}(P^{-1}M^{\dagger}Q^{-1}M)-{\rm Tr}\ln({\mathbb{1}}^{\otimes 2}-iL\otimes(KM^{\dagger}M))}\;.$ (5.30)
As in Section 5.1, contractions of the first index of the matrix $M$ generate color–1 faces weighted by ${\rm Tr}(Q^{n})$. Given the form of the dually weighted potential $V_{\ln}[K,L]$, each vertex carries a weight ${\rm Tr}(L^{n})$, just as in the self-adjoint one-matrix model. However, in this case, the face structure is enriched by the presence of two independent families of cycles built into the complex matrix model. This is because contractions of the second index of the matrix $M$ generate color–2 faces weighted by ${\rm Tr}((PK)^{n})$. In summary, for each $n$, one obtains
@@
$\mathcal{Z}_{\rm HM}[V_{\ln}[K,L],Y_{\ln}[P,Q]]=\int dAdB\;e^{-iN{\rm Tr}(AB)-{\rm Tr}\ln({\mathbb{1}}^{\otimes 2}-iL\otimes(KA))-{\rm Tr}\ln({\mathbb{1}}^{\otimes 2}-iQ\otimes(PB))}\;.$ (5.31)
@@
  • AA-vertices of degree $n$ each with weight ${\rm Tr}(L^{n})$.
  • $B$-vertices of degree $n$ each with weight ${\rm Tr}(Q^{n})$.
  • Faces of length $n$ each with weight ${\rm Tr}((PK)^{n})$.
This is a bipartite-graph version of dually weighted matrix models [26, 13], and according to Theorem 5.3, we have
$\frac{\mathcal{Z}_{\rm CM}[V_{\ln}[K,L]](P,Q)}{\mathcal{Z}_{\rm CM}[0](P,Q)}=\frac{\mathcal{Z}_{\rm HM}[V_{\ln}[K,L],Y_{\ln}[P,Q]]}{\mathcal{Z}_{\rm HM}[0,Y_{\ln}[P,Q]]}\;.$ (5.32)
@@
An interesting case arises when $Q=C_{2}$ (See (3.6)). We recall that the matrix $C_{2}$ is only well defined in the large $N$ limit [7]. Therefore, consider the large-$N$ partition function [^9] This is similar to the partition function seen in (5.12), but it uses the matrix $C_{2}$, which is only well-defined in the large-N limit.
$\mathcal{Z}_{\rm HM}[V,Y_{\ln}[P,C_{2}]]=\int dAdB\;e^{-iN{\rm Tr}(AB)+\frac{1}{2}N{\rm Tr}((PB)^{2})+V(A)}\;.$ (5.33)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,C_{2}]]=N^{-n}\sum_{\gamma\in S_{n}}{\rm Tr}_{[\gamma\sigma]}(P){\rm Tr}_{[\gamma]}(C_{2})\;.$ (5.35)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,C_{2}]]=0\qquad\text{if $n$ is odd}\;,$ (5.36)
@@
$\left\langle{\rm Tr}_{[\sigma]}(A)\right\rangle[0,Y_{\ln}[P,C_{2}]]=N^{-m}\sum_{\gamma\in[2^{m}]}{\rm Tr}_{[\gamma\sigma]}(P)\qquad\text{if $n=2m$ is even}\;.$ (5.37)
@@
for $n=2m$ even and $1\leq n\leq N$. This correspondence is illustrated in Figure 11.
Figure 11: An example of equivalence between a graph in the: (a) complex matrix theory ${\cal Z}_{\rm CM}[V](P,C_{2})$ (5.38); (b) intermediate theory (5.21); (c) the Hermitian matrix theory $\mathcal{Z}_{\rm HM}[V,Y_{\ln}[P,C_{2}]]$ (5.15); (d) the Hermitian matrix theory $\mathcal{Z}_{\rm HM}[V,Y_{\ln}[P,C_{2}]]$ (5.34) after integrating $B$. The dashed blue lines represent the propagator of $A$ in (5.11). Intermediate fields AA and $B$ are represented in gray lines.
@@
where ${\rm Cat}(m)$ are the Catalan numbers. If $V(M^{\dagger}M)=-N\frac{g^{2}}{2}(M^{\dagger}M)^{2}$, then
$\mathcal{Z}_{\rm HM}[V,Y_{\ln}[{\mathds{1}},C_{2}]]=\int dA\;e^{-N{\rm Tr}\left(\frac{1}{2}A^{2}-\frac{g^{2}}{2}A^{2}\right)}\;,$ (5.41)
showing that the model (3.10) becomes equivalent to the Gaussian model above (5.41). Appendix E.1 explores the case when $Q=C_{2}$ further via explicit graph resummation.
$\mathbf{Q=C_{k}}$: $k$-uniform rigidity and a partition-matroid structure.
@@
$\big\langle{\rm Tr}_{[\sigma]}(M^{\dagger}M)\big\rangle[0](P,C_{k})=\big\langle{\rm Tr}_{[\sigma]}(A)\big\rangle[0,Y_{\ln}[P,C_{k}]]=N^{-d}\sum_{\gamma\in[k^{d}]}{\rm Tr}_{[\gamma\sigma]}(P).$ (5.45)
@@ 5.5 Miscellaneous results
Let $A\in\mathrm{Herm}(N)$ be a random matrix distributed according to ${\cal Z}_{\mathrm{HM}}[0,Y_{\ln}[P,Q]]$. Then, for any representation $r$ of ${\rm GL}(N)$,
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=n!N^{-n}\frac{1}{\chi_{r}(e)}\chi_{r}(P)\chi_{r}(Q)\;,$ (5.47)
@@
$\chi_{r}(A)=\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi_{r}(\sigma){\rm Tr}_{[\sigma]}(A)\;.$ (5.48)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=\frac{1}{n!}N^{-n}\sum_{\begin{subarray}{c}\sigma,\gamma\in{\rm S}_{n}\end{subarray}}\chi_{r}(\sigma){\rm Tr}_{[\gamma\sigma]}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (5.49)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=\frac{1}{n!}N^{-n}\sum_{\begin{subarray}{c}\sigma,\gamma\in{\rm S}_{n}\end{subarray}}\chi_{r}(\sigma)\sum_{s}\chi_{s}(\gamma\sigma)\chi_{s}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (5.50)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=\frac{1}{n!}N^{-n}\sum_{\begin{subarray}{c}\sigma,\gamma\in{\rm S}_{n}\end{subarray}}\chi_{r}(\sigma)\sum_{s}\frac{\chi_{s}(\gamma)\chi_{s}(\sigma)}{\chi_{s}(e)}\chi_{s}(P){\rm Tr}_{[\gamma]}(Q)\;,$ (5.51)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=N^{-n}\sum_{\begin{subarray}{c}\gamma\in{\rm S}_{n}\end{subarray}}\sum_{s}\frac{\chi_{s}(\gamma)\delta_{rs}}{\chi_{s}(e)}\chi_{s}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (5.52)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=N^{-n}\sum_{\begin{subarray}{c}\gamma\in{\rm S}_{n}\end{subarray}}\frac{\chi_{r}(\gamma)}{\chi_{r}(e)}\chi_{r}(P){\rm Tr}_{[\gamma]}(Q)\;.$ (5.53)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=n!N^{-n}\frac{1}{\chi_{r}(e)}\chi_{r}(P)\chi_{r}(Q)\;.$ (5.54)
This expression is valid for any $N>0$ and $n\geq 0$. If $n\leq N$, we can use $\chi_{r}(e)=N^{-n}n!\chi_{r}(C_{1})$. For the following corollary, we recall the matrix $C_{1}$ given in (3.6), the partition function ${\cal Z}_{\mathrm{HM}}$ given in (5.7), and the expectation value $\langle\cdot\rangle$ given in (5.8).
@@
Let $A\in\mathrm{Herm}(N)$ be a random matrix distributed according to ${\cal Z}_{\mathrm{HM}}[0,Y_{\ln}[P,Q]]$. Then, for any representation $r$ of ${\rm GL}(N)$ with size $n\leq N$,
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=\frac{\chi_{r}(P)\chi_{r}(Q)}{\chi_{r}(C_{1})}\;.$ (5.55)
Since the character $\chi_{r}(e)$ does not depend on $N$, one has that, at fixed $n$,
$\chi_{r}(e)=n!\lim_{N\rightarrow\infty}N^{-n}\chi_{r}(C_{1})\;.$ (5.56)
@@
$\left\langle\chi_{r}(A)\right\rangle[0,Y_{\ln}[P,Q]]=\frac{\chi_{r}(P)\chi_{r}(Q)}{\tilde{\chi}_{r}(C_{1})}\;.$ (5.58)
@@
$\left\langle f(A)\right\rangle[0,Y_{\ln}[P,Q]]=\sum_{n=0}^{\infty}\sum_{r\vdash n}f^{r}\frac{\chi_{r}(P)\chi_{r}(Q)}{\tilde{\chi}_{r}(C_{1})}\;.$ (5.60)
@@
$\chi_{r}(M^{\dagger}M)=\frac{1}{n!}\sum_{\sigma\in{\rm S}_{n}}\chi_{r}(\sigma){\rm Tr}_{[\sigma]}(M^{\dagger}M)\;,$ (5.61)
@@ 6.1 The complex tensor model
If $\varphi$ is a random tensor distributed according to the partition function ${\cal Z}_{\rm CT}[0](R)$ (6.1), then, for any $\bm{\sigma}\in{\rm S}_{n}^{D}$, the $\bm{\sigma}$}–bubble/trace invariant satisfies
@@
$\langle\varphi\varphi^{\dagger}\rangle[0](R)=\bigotimes_{c=1}^{D}R_{c}$ (6.6)
and is illustrated in Figure 12.
Figure 12: Ribbon graph representation of the propagator of the theory defined by the partition function (6.1) as given in (6.6). The blue, green, and red strands respectively represent the presence of $R_{1},R_{2}$, and $R_{3}$ in $R=R_{1}\otimes R_{2}\otimes R_{3}$.
Furthermore, the expectation value (6.4) factorizes accordingly, as stated in the following Corollary.
@@
This factorized expression has a clear geometric meaning in the stranded Feynman-graph representation of tensor models: it decomposes into $D$ independent families of faces, one for each tensor index. This is because when the covariance factorizes as $R=\bigotimes_{c=1}^{D}R_{c}$, the tensor propagator splits into $D$ independent strands, where the $c$-th strand represents the propagator weight $R_{c}$. This implies that the Wick contractions separate by color: the same permutation $\mu\in S_{n}$ governs the pairings of all tensor slots, but in each color $c$, it combines with the bubble permutation $\sigma_{c}$ to form $\mu\sigma_{c}$. The cycles of $\mu\sigma_{c}$ are precisely the faces traced by the $c$-th index of the tensor, and a cycle of length $l$ corresponds to a color-$c$ face of length $l$ in the stranded graph. Since a color-$c$ strand represents a propagator with weight $R_{c}$, a color-$c$ face of length $l$ has a weight ${\rm Tr}(R_{c}^{l})$. This is the tensor analogue of the dually weighted structure of the complex matrix model presented in Section 5.1, now extended to $D$ colors.
@@
In particular, the scalar-$U(1)$ and partial-$U(N)$ cases will be relevant for the equivalence theorems we prove, and we elaborate on these below. Notice that scalar-$U(1)$ $\subset$ partial-$U(N)$ $\subset$ full-$U(N)^{D}$, and that scalar-$U(1)$ $\subset$ repeated-$U(N)$ $\subset$ full-$U(N)^{D}$, but there is no inclusion relation between partial-$U(N)$ and repeated-$U(N)$.
@@
The difference between a general $W(\varphi,\varphi^{\dagger})$ and a potential of the form $V(\varphi\varphi^{\dagger})$ is illustrated in Figure 13: in the former, $\varphi$ and $\varphi^{\dagger}$ may appear in arbitrary orderings or in different ammounts, while in the latter they are naturally paired into copies of $\varphi\varphi^{\dagger}$.
Figure 13: We represent a few examples of potentials via stranded notation in the case $D=3$. Some examples include a constant tensor $\psi$ independent of $\varphi$. Some strands may be colored if they exhibit a partial-$U(N)$ symmetry; the first, second and third indexes of $\varphi$ are respectively represented in red, green and blue. The examples are: (a) A potential $W$ without scalar-$U(1)$ symmetry. (b) A potential $V$ with scalar-$U(1)$ symmetry but without repeated-$U(N)$ symmetry. (c) A potential $V$ with repeated-$U(N)$ symmetry but without partial-$U(N)$ symmetry. (d) A potential $V$ with partial-$U(N)$ symmetry but without full-$U(N)^{3}$ symmetry. (e) A potential $V$ with full-$U(N)^{3}$ symmetry.